# MM-Interleaved: Interleaved Image-Text Generative Modeling via Multi-modal Feature Synchronizer

### v1

@@ 3.2 Architecture
Building upon the task formulation in Sec. 3.1, we propose a novel multi-modal architecture for processing interleaved image-text data. It integrates a Visual Foundation Model (VFM), a Large Language Model (LLM), and a Diffusion Model (DM). This integration aims to excel in both understanding and generation tasks of text and images by leveraging the strengths of each model type. As illustrated in Fig. 3, our architecture comprises three key components:
(1) VFM-based Image Tokenizer $\mathcal{E}_{V}$ that maps each image $x^{V}\in\mathbb{R}^{H\times W\times 3}$ (e.g., $H=W=224$) into a fixed number of visual tokens $e^{V}\in\mathbb{R}^{N\times C}$ ($N=32$ by default). $C$ is the channel dimension. It consists of a pre-trained vision foundation model (e.g., CLIP-ViT [70]) for feature extraction and a Perceiver Resampler [3] to reduce the number of visual tokens. We also use a ViT-Adapter [12] to extract multi-scale image features $F^{V}\in\mathbb{R}^{(\sum_{i=1}^{L}H_{i}\times W_{i})\times C}$ for fine-grained feature fusion in subsequent networks, where $L=4$ and $H_{i}=\frac{H}{2^{i+1}},W_{i}=\frac{W}{2^{i+1}}$ by default.
(2) LLM-based Multi-modal Model $\mathcal{D}_{\text{LM}}$ that extracts context features from the interleaved image-text sequences. A pre-trained LLM (e.g., Vicuna [105]) is utilized. Its input sequence $E\in\mathbb{R}^{K\times C}$ is a concanation of embeddings $(e_{1},e_{2},\dots)$, where $e_{n}$ is either a word embedding $e_{n}^{L}\in\mathbb{R}^{1\times C}$ or an image embedding $e_{n}^{V}\in\mathbb{R}^{N\times C}$. $K$ is the total number of input tokens. We also introduce a fine-grained feature synchronizer module for letting the intermediate layers in $\mathcal{D}_{\text{LM}}$ directly access and extract multi-scale image features on demand.
@@
Figure 4: Architecture of the MMFS module. The query feature is passed through a linear layer and added to the image index embedding. Two linear layers are used to predict the sampling offsets and unnormalized attention weights of each image respectively. The sampling offsets are added with the query’s reference point to form the corresponding sampling locations, which are shared and rescaled across all feature maps of the same image. The output feature is the weighted sum of the sampled spatial features from these points.
Multi-image and Multi-scale Feature Synchronizer (MMFS). The MMFS module is designed to dynamically capture image details from multiple images in an interleaved context. It leverages the Deformable Attention [110] mechanism to achieve efficient and sparse image attention. As is shown in Fig. 4, given a query token that requires detailed image features, the MMFS module only attends to a small set of sampling points around a reference point on the reference images. Let $f_{q}\in\mathbb{R}^{C}$ represent the features of the query token. $\hat{p}_{q}\in[0,1]^{2}$ denotes the relative image coordinate of its reference point. By default, $\hat{p}_{q}=(0.5,0.5)$ corresponds to the center of the images. $\{F^{V}_{m}\}_{m=1}^{M}$ are the multi-image multi-scale feature maps, where $F^{V}$ is the multi-scale feature extracted by the image tokenizer, $M$ is the number of reference images. The output feature $f_{o}\in\mathbb{R}^{C}$ is given by
@@
$p_{q}$ $=\mathrm{Concat}(p_{q}^{(1)},\cdots,p_{q}^{(M)}).$
$A_{q}$ $=\mathrm{softmax}\left(\mathrm{Concat}(A_{q}^{(1)},\cdots,A_{q}^{(M)})\right).$
@@
MMFS can be applied to both image and text decoding, avoiding information bottlenecks caused by the widely used Resamplers in multi-modal LLMs. It is especially efficient for processing multiple high-resolution images in context.
Multi-modal LLM with MMFS. The input of multi-modal LLMs is an interleaved sequence of image and text token embeddings, which always starts with a special token <s> and ends with another special token </s>. Image token embeddings are inserted at the corresponding positions in the original sequence. Another special token <BoI> is added in front of each image to represent “Begin of Image”.
MMFS modules are inserted between the self-attention layer and the feed-forward layer of the LLM every fixed number of blocks. Query token $f_{q}$ iterates over each token in the LLM, which can only access the previous images. $\hat{p}_{q}=(0.5,0.5)$. The output of MMFS is multiplied by tanh($\alpha$) before being added back to $f_{q}$, where $\alpha$ is a zero-initialized learnable scalar.
Image Decoder with MMFS. The output features from the multi-modal LLM are further processed by another Perceiver Resampler before being fed into the image decoder. For each image to be generated, the Resampler maps previous context features to a fixed number of tokens (e.g., 77 tokens) to match the pre-trained diffusion model.
MMFS modules are inserted after each upsampling block in the U-Net structure of the diffusion model. Query token $f_{q}$ iterates over each pixel in the feature map. $\hat{p}_{q}$ is set as the spatial coordinate of the query pixel. The output of MMFS is further processed by a zero-initialized convolution layer before being added back to $f_{q}$.
Training Target and Inference Pipeline. The training objective is defined as the sum of Next-Text-Token Prediction loss in Eq. (2) and Next-Image Prediction loss in Eq. (3),
$\mathcal{L}=\mathcal{L}_{NTP}+\lambda~{}\mathcal{L}_{NIP}$ (5)
where $\lambda$ is a hyperparameter used to determine the relative loss weight between the image and text decoding branches. The whole framework can be optimized end-to-end.
@@ 4.1 Implementation Details
[…]
@@
MM-Interleaved Vicuna-13B - - - 129.0 85.8 106.4 23.5 57.0 40.0 40.8 37.2 48.7
@@
[…]
Network. We adopt CLIP-ViT-L/14 [70], Vicuna-13B [105] and Stable Diffusion v2.1 [74] as the image encoder, large language model, and image decoder, respectively. For the multi-modal LLM, a Perceiver Resampler with 12 blocks is used to reduce the number of visual tokens per image to 64. The MMFS module is inserted every 4 blocks in the LLM. For the image decoder, a Perceiver Resampler with only 1 block is used to reduce the number of conditional tokens per image to 77. The MMFS module is inserted after each up-sampling block in the image decoder.
Pre-training. Our models are pre-trained on a mixture of interleaved image-text sequences and image-text pairs, including MMC4 [109], LAION-2B [76], LAION-COCO [77], CC-12M [8] and Objects365 [78]. For CC-12M [8] and Objects365 [78], instead of utilizing the original annotations, we use the pre-trained BLIP-2 model [48] to caption the images. The sampling probability of MMC4 is twice that of other image-text pair datasets. No in-house data is used.
The images are inserted before or after the corresponding text sentence with equal probability. To optimize training efficiency and data utility, multiple image-text pairs or interleaved image-text sequences are concatenated into extended sequences with the maximum context length (i.e., 2,048 tokens).
The model is pre-trained for 15,000 steps with 11 billion tokens visited. The image encoder and LLM are frozen. The learning rate is set to be $10^{-5}$ for the image decoder and $10^{-4}$ for the rest trainable parameters. The input and output image resolutions are $224\times 224$ and $512\times 512$, respectively.
Supervised Fine-tuning. After pre-training, like multi-modal LLMs, our model capabilities can be further enhanced using supervised fine-tuning. The model is fine-tuned on four types of downstream tasks, which are 1) visual question-answering and image caption, 2) referring expression comprehension, 3) segmentation-to-image translation, and 4) visual storytelling.
More details of datasets and hyper-parameters used for pre-training and fine-tuning could be found in the Appendix.
Model RefCOCO [42] RefCOCO+ [58] RefCOCOg [58]
Val Test-A Test-B Val Test-A Test-B Val Test
OFA-L [88] 79.96 83.67 76.39 68.29 76.00 61.75 67.57 67.50
VisionLLM-H [90] - 86.70 - - - - - -
Shikra [10] 87.01 90.61 80.24 81.60 87.36 72.12 82.27 82.19
MiniGPT-V2 [9] 88.69 91.65 85.33 79.97 85.12 74.45 84.44 84.66
Ferret [96] 89.48 92.41 84.36 82.81 88.14 75.17 85.83 86.34
 Qwen-VL [6] 89.36 92.26 85.34 83.12 88.25 77.21 85.58 85.48
MM-Interleaved (ours) 89.92 92.59 86.54 82.99 88.57 77.07 85.21 84.92
Table 2: Supervised fine-tuning results on referring expression comprehension task. “*” denotes using an additional self-constructed grounding dataset and trained with an image resolution larger than 224.
@@
MM-Interleaved-13B (ours) 7.90 23.88
Table 3: Zero-shot text-to-image generation results on MS-COCO [51] and LN-COCO [69]. FID [31] is reported.
@@ 4.2 Evaluation
We evaluate the zero-shot visual-language capability of the pre-trained MM-Interleaved as well as its performance on various downstream tasks after fine-tuning and report the results in Tab. 1.
@@
Zero-shot Results after Pre-training. Our method demonstrates strong multi-modal zero-shot comprehension on various benchmarks, including image captioning on COCO [11], Flicker30K [68], NoCaps [2] and Image2Paragraph [44]; general visual question answering on VQAv2 [27], OKVQA [59], VizWiz [29] and TextVQA [80]. We also evaluate on the VisDial [15] visual dialogue task. In the fully decontaminated setting where the image and text in downstream tasks are unseen during pre-training, our model significantly outperforms MetaLM [30], OF-9B [5] and IDEFICS-80B [45] on all captioning, VQA, and multi-modal tasks. This demonstrates the effectiveness of the proposed MM-Interleaved approach. Furthermore, our model even exceeds most trained on vast in-house data like Flamingo-9B [3] and Emu [83], showing the importance of proper architecture for image-text interaction.
@@
Additionally, we evaluate text-conditional image generation on MS-COCO [51] and LN-COCO [69]. On MS-COCO, we sample 8 images per text condition and use CLIP [70] to rerank based on text-image similarity. CLIP reranking is not used for LN-COCO. FID [31] is used to evaluate both datasets. As show in Tab. 3, our model shows competitive text-to-image generation compared to existing image and text generation models.
@@
Fine-tuning Results of VQA and Captioning. Our fine-tuned model achieves state-of-the-art performance on image captioning benchmarks. On visual question answering tasks, it matches the previous best LLaVA-1.5 model [52]. Compared to LLaVA-1.5, our model has two key advantages 1) our method is capable of generating both images and text, while LLaVA-1.5 can only generate text; 2) LLaVA-1.5 uses 576 visual tokens as LM inputs, whereas we only require 64. Our model achieves competitive image understanding with far fewer visual tokens, making it better suited for multi-image scenarios.
@@
Fine-tuning Results of Referring Expression Comprehension. Our model demonstrates state-of-the-art performance on various referring expression comprehension (REC) benchmarks, as shown in Tab. 2. Compared to recent REC models like VisionLLM [90] and Shikra [10], our model outperforms on most datasets. Remarkably, even though we only use public REC data [42, 58] for fine-tuning, our model matches QWen-VL [6] which utilizes a extra 22M in-house grounding dataset and trains at higher resolution (448 vs 224 pixels). This shows the effectiveness of synchronized fine-grained image features for enhancing the REC capability of our model.
@@
Fine-tuning Results of Segmentation-to-Image Translation. We follow the protocol proposed in ControlNet [100] by generating images from ADE20K ground truth semantic masks and use OneFormer [38] for segmenting the generated images. Standard mIoU is reported as a metric for measuring the semantic and structural fidelity of the generated images. We show the results in Tab. 3(a). MM-Interleaved clearly outperforms other baselines including ControlNet by a large margin. Compared to ControlNet, our model could leverage the better representations learned by our generative modeling from large-scale pre-training data, which benefits the image generation tasks. Moreover, thanks to MMFS, our model is capable of generating realistic images with pixel-level precise alignment from a semantic label map, where simply relying on coarse Perceiver Resampler features fails on this task, which we will show in the ablation studies.
ADE20K (GT) [38] VQGAN [20] LDM [73]
@@
(a) Segmentation-to-image generation results on ADE20K [106] mIoU ($\uparrow$) is reported.
@@
(b) Image generation results on VIST [34].
@@
(c) Autoregressive image generation results on Pororo [49] and FlintStones [28]. FIDs($\downarrow$) are reported.
Table 4: Comparative analysis of image generation ability of MM-Interleaved.
Fine-tuning Results of Visual Storytelling. Besides text-to-image and image-to-image generation, our MM-Interleaved could also condition on interleaved image-text as context for generating new images. We evaluate our model’s performance both for generating the last image on the Visual StoryTelling (VIST) dataset [34] and for generating multiple images autoregressively on the Pororo [49] and Flintstones [28] datasets. As is shown in Tab. 3(b) and  3(c), our model achieves SoTA performance on both CLIP-Similarity and FID metrics compared to previous works on VIST and achieves competitive performance compared with previous specialist methods on Pororo and Flinestones. More qualitative results and implementation details can be found in the appendix.

### v2

@@ 3.2 Architecture
Building upon the task formulation in Sec. 3.1, we propose a novel multi-modal architecture for processing interleaved image-text data. It integrates a Visual Foundation Model (VFM), a Large Language Model (LLM), and a Diffusion Model (DM). Such integration aims to excel in both comprehension and generation tasks of text and images by leveraging the strengths of each model type. As illustrated in Fig. 4, our architecture comprises three key components:
(1) VFM-based Image Tokenizer $\mathcal{E}_{V}$ that maps each image $x^{V}\in\mathbb{R}^{H\times W\times 3}$ (e.g., $H=W=224$) into a fixed number of visual tokens $e^{V}\in\mathbb{R}^{N\times C}$ ($N=32$ by default). $C$ is the channel dimension. It consists of a pre-trained vision encoder (e.g., CLIP-ViT [71]) for feature extraction and a Perceiver Resampler [3] to reduce the number of visual tokens. A ViT-Adapter [12] is also used to extract multi-scale image features $F^{V}\in\mathbb{R}^{(\sum_{i=1}^{L}H_{i}\times W_{i})\times C}$ for fine-grained feature fusion in subsequent networks, where $L=3$ and $H_{i}=\frac{H}{2^{i+2}},W_{i}=\frac{W}{2^{i+2}}$ by default.
(2) LLM-based Multi-modal Model $\mathcal{D}_{\text{LLM}}$ that extracts context features from the interleaved image-text sequences. A pre-trained LLM (e.g., Vicuna [107]) is utilized. Its input sequence $E\in\mathbb{R}^{K\times C}$ is a concatenation of embeddings $(e_{1},e_{2},\dots)$, where $e_{n}$ is either a word embedding $e_{n}^{L}\in\mathbb{R}^{1\times C}$ or an image embedding $e_{n}^{V}\in\mathbb{R}^{N\times C}$. $K$ is the total number of input tokens. We also introduce a feature synchronizer (i.e., MMFS) for letting the intermediate layers in $\mathcal{D}_{\text{LLM}}$ directly access and extract multi-scale image features on demand.
@@
Multi-Modal Feature Synchronizer (MMFS). MMFS aims to enable dynamic and efficient image detail extraction in decoder intermediate layers, compensating for information loss due to limited input visual tokens. It leverages the Deformable Attention [112] to achieve efficient and sparse image attention. MMFS can be applied to both image and text decoding, avoiding information bottlenecks caused by the Resamplers in multi-modal LLMs. It is especially efficient for processing multiple high-resolution images in context.
Figure 5: Architecture of MMFS module. The query feature is passed through a linear projection and added with the image index embedding. Two linear projections are used to predict the sampling offsets and unnormalized attention weights of each image, respectively. The sampling offsets are added with the query’s reference point to form the corresponding sampling locations, which are shared across all feature maps of the same image. The output is the weighted sum of the sampled spatial features.
As is shown in Fig. 5, given a query token that requires detailed image features, the MMFS module only attends to a small set of sampling points around a reference point on the reference images. Let $f_{q}\in\mathbb{R}^{C}$ represents the features of the query token. $\hat{p}_{q}\in[0,1]^{2}$ denotes the relative image coordinate of its reference point, which is designed for compressing the search space for sampling points from global positions to local relative offsets. By default, without a spatial prior, $\hat{p}_{q}=(0.5,0.5)$ is set as the center of the images. $\{F^{V}_{m}\}_{m=1}^{M}$ are the multi-image multi-scale feature maps extracted by the image tokenizer, $M$ is the number of reference images. The output feature $f_{o}\in\mathbb{R}^{C}$ is given by
@@
$p_{q}$ $=\mathrm{Concat}(p_{q}^{(1)},\cdots,p_{q}^{(M)}),$
$A_{q}$ $=\mathrm{softmax}\left(\mathrm{Concat}(A_{q}^{(1)},\cdots,A_{q}^{(M)})\right),$
@@
Multi-modal LLM with MMFS. The input of multi-modal LLMs is an interleaved sequence of image and text token embeddings, which always starts with a special token <s> and ends with another special token </s>. Image token embeddings are inserted at the corresponding positions in the original sequence. Special token <BoI> is added in front of each image to represent “Begin of Image”.
MMFS modules are inserted between the self-attention layer and the feed-forward layer of the LLM every fixed number of blocks. Query token $f_{q}$ iterates over each token in the LLM, which can only access the previous images. $\hat{p}_{q}=(0.5,0.5)$. The output of MMFS is multiplied by tanh($\alpha$) before added back to $f_{q}$, where $\alpha$ is a zero-initialized learnable scalar.
Image Decoder with MMFS. The output features from the multi-modal LLM are further processed by another Resampler before fed into the image decoder. For each image to be generated, the Resampler maps previous context features to a fixed number of tokens (e.g., 77 tokens) to match the condition token length of the pre-trained diffusion model.
MMFS modules are inserted after each downsampling block in the U-Net of the diffusion model. Query token $f_{q}$ iterates over each pixel in the feature map. $\hat{p}_{q}$ is set as the spatial coordinate of the query pixel. The output of MMFS is further processed by a zero-initialized convolution before added back to $f_{q}$.
Training Target and Inference Pipeline. The training objective is defined as the sum of Next-Text-Token Prediction loss in Eq. (2) and Next-Image Prediction loss in Eq. (3) as $\mathcal{L}=\mathcal{L}_{NTP}+\lambda~{}\mathcal{L}_{NIP}$, where $\lambda$ is a hyperparameter used to determine the relative loss weight between the image and text decoding branches. The whole framework can be optimized end-to-end.
@@ 4.1 Implementation Details
Network. We adopt CLIP-ViT-L/14 [71], Vicuna-13B [107] and Stable Diffusion v2.1 [74] as the image encoder, large language model, and image decoder, respectively. For the multi-modal LLM, a Perceiver Resampler with 12 blocks is used to reduce the number of visual tokens per image to 64. MMFS is inserted every 4 blocks in the LLM. For the image decoder, a Perceiver Resampler with only 1 block is used to reduce the number of previous conditional tokens to 77. MMFS is inserted after each downsampling block in the image decoder.
Pre-training. Our model is pre-trained on a mixture of image-text pairs and interleaved image-text sequences, including MMC4 [111], LAION-2B [77], LAION-COCO [78], CC-12M [8] and Objects365 [79]. For CC-12M [8] and Objects365 [79], instead of utilizing the original annotations, we use the pre-trained BLIP-2 model [47] to caption the images. The sampling probability of MMC4 is twice that of other image-text pair datasets. No in-house data is used.
The images are inserted before or after the corresponding text sentence with equal probability. To optimize training efficiency and data utility, multiple image-text pairs or interleaved image-text sequences are concatenated into extended sequences with the maximum context length (i.e., 2048 tokens).
The model is pre-trained for 15000 steps with 11 billion tokens visited. The image encoder and LLM are frozen. The learning rate is set to be $10^{-5}$ for the image decoder and $10^{-4}$ for the rest trainable parameters. The input and output image resolutions are $224\times 224$ and $512\times 512$, respectively.
Model LLM H I A COCO Flickr NoCaps I2Para. VQAv2 OKVQA GQA VizWiz TextVQA VisDial
@@
VL-GPT-I [110] LLaMA-7B - I A 133.7 67.2 50.3 51.5 38.9 51.8
SEED-LLaMA [25] LLaMA2-Chat-13B - I A 125.0 48.1 27.1 23.3
SEED-LLaMA-I [25] LLaMA2-Chat-13B - I A 126.9 63.4 43.2 49.4
MM-Interleaved Vicuna-13B - - - 129.0 85.8 106.4 23.5 57.0 40.0 46.7 40.8 37.2 48.7
@@
Table 1: Multi-modal comprehension evaluation. “H” denotes using in-house data, “I” means the training images of some benchmarks are included in the training, “A” means the training annotations of some benchmarks are visible in training. Benchmarks include COCO [11]; Flickr: Flickr30k [69]; NoCaps [2]; I2Para.: Image2Paragraph [44]; VQAv2: VQAv2 [27]; OKVQA [60]; GQA [35]; VizWiz [29]; TextVQA [81]; VisDial [15]. The results in underline and bold are the best performance without “HIA” data and without “H” data, respectively. The evaluation metrics for each benchmark are listed in Appendix.
Model LLM $L_{1}$ (Part-1) Part-2 $L_{2}$ Part-3 $L_{3}$
Emu [84] LLaMA-13B 42.5 41.1 42.4 41.4 42.3
Next-GPT [95] Vicuna-7B 30.7 35.6 31.1 33.9 31.4
SEED-LLaMA [25] LLaMA2-Chat-13B 43.9 43.4 43.8 52.3 44.8
MM-Interleaved Vicuna-13B 43.9 46.1 44.1 52.1 45.0
Table 2: Zero-shot results for interleaved image-text comprehension and generation on SEED-Bench-2 [46]. The average task accuracy across corresponding evaluation dimensions is reported. $L_{1}$(part-1) evaluates the image and text comprehension, $L_{2}$(part-1&2) evaluates interleaved image-text comprehension, and $L_{3}$(part-1&2&3) evaluates image and text generation.
@@
MM-Interleaved 7.90 23.88
Table 3: Zero-shot text-to-image generation results. FID [31] is reported.
Supervised Fine-tuning. After pre-training, like other multi-modal LLMs, our model capabilities can be further enhanced using supervised fine-tuning. The model is end-to-end fine-tuned on four types of downstream tasks: 1) visual question-answering and image caption, 2) visual storytelling, 3) segmentation-to-image translation, and 4) referring expression comprehension.
More implementation details could be found in Appendix.
@@ 4.2 Evaluation
We evaluate the zero-shot capability of the pre-trained MM-Interleaved as well as its performance on various downstream tasks after supervised fine-tuning.
@@
Zero-shot Results after Pre-training. Tab. 1 demonstrates our strong performance on multi-modal zero-shot comprehension across various benchmarks, including image captioning (COCO [11], Flicker30K [69], NoCaps [2], I2Para [44]), visual question answering (VQAv2 [27], OKVQA [60], VizWiz [29], TextVQA [81]) and visual dialogue (VisDial [15]). In the fully decontaminated setting where the image and text in downstream tasks are unseen during pre-training, our model significantly outperforms other methods on all tasks. This demonstrates the effectiveness of the proposed MM-Interleaved approach. Furthermore, our model even exceeds most trained on vast in-house data like Flamingo-9B [3] and Emu [84], showing the importance of proper architecture for image-text interaction.
@@
Tab. 3 shows the results on text-to-image generation for MS-COCO [52] and LN-COCO [70]. On MS-COCO, we sample 8 images per text condition and use CLIP [71] to rerank based on text-image similarity. CLIP reranking is not used for LN-COCO. Our model achieves competitive performance compared to existing image and text generation models. Note that some other works (e.g., Emu, Muse, Imagen, and Parti) use in-house data, while ours did not.
Tab. 2 shows the zero-shot performance for interleaved image-text comprehension and generation on SEED-Bench-2 [46]. Our pre-trained model achieves SOTA results on both comprehension and generation tasks. Visualization results for interleaved image-text generation are showed in Appendix.
Model RefCOCO [42] RefCOCO+ [59] RefCOCOg [59]
Val Test-A Test-B Val Test-A Test-B Val Test
OFA-L [89] 79.96 83.67 76.39 68.29 76.00 61.75 67.57 67.50
VisionLLM-H [92] - 86.70 - - - - - -
Shikra [10] 87.01 90.61 80.24 81.60 87.36 72.12 82.27 82.19
MiniGPT-V2 [9] 88.69 91.65 85.33 79.97 85.12 74.45 84.44 84.66
Ferret [98] 89.48 92.41 84.36 82.81 88.14 75.17 85.83 86.34
Qwen-VL [6] 89.36 92.26 85.34 83.12 88.25 77.21 85.58 85.48
MM-Interleaved 89.92 92.59 86.54 82.99 88.57 77.07 85.21 84.92
Table 4: Supervised fine-tuning results on referring expression comprehension task. * denotes using an additional self-constructed grounding dataset and trained with an image resolution larger than 224.
@@
Fine-tuning Results. Tab. 1 also shows our fine-tuned results (MM-Interleaved-SFT) on multi-modal comprehension. Our fine-tuned model achieves SOTA performance without using any in-house data. On visual question answering tasks, it matches the previous best LLaVA-1.5 model. Compared to LLaVA-1.5, our model has two key advantages: 1) our method is capable of generating both images and text, while LLaVA-1.5 can only generate text; 2) LLaVA-1.5 uses 576 visual tokens as LLM inputs, whereas we only require 64 tokens. Our model achieves competitive image understanding with far fewer visual tokens, making it better suited for multi-image scenarios.
@@
Tab. 4 shows the results on referring expression comprehension (REC) benchmarks. Our method outperforms other methods. Even though we only use public REC data [42, 59] for fine-tuning, MM-Interleaved matches Qwen-VL [6] which utilizes an extra 22M in-house grounding dataset and trains at higher resolution (448 vs. 224 pixels used by ours). This shows the effectiveness of synchronized fine-grained image features for enhancing the REC capability of our method.
@@
Tab. 5(a) shows the performance on segmentation-to-image translation, which is a capability not shown in other multi-modal LLMs. We follow the protocol proposed in ControlNet [102] by generating images from ADE20K ground truth semantic masks and use OneFormer [38] for segmenting the generated images. Standard mIoU metric is used for measuring the semantic and structural fidelity of the generated images. MM-Interleaved clearly outperforms other baselines including ControlNet by a large margin. Compared to ControlNet, MM-Interleaved could leverage the better representations learned by our generative modeling from large-scale pre-training data, which benefits the image translation tasks. Moreover, thanks to MMFS, MM-Interleaved is capable of generating realistic images with pixel-level precise alignment from a semantic label map, where simply relying on coarse Perceiver Resampler features fails on this task (see Tab. 6(c)).
Groundtruth VQGAN [19] LDM [75]
@@
(a) Segmentation-to-image generation on ADE20K [108]. mIoU is reported.
@@
(b) Last image generation with interleaved context for visual storytelling on VIST [34].
@@
(c) Multi-image generation with interleaved context for visual storytelling. FID is reported.
Table 5: Segmentation-to-image generation and visual storytelling results.
Tabs. 5(b) and 5(c) show our results on visual storytelling, which is required to generate new images with interleaved image-text as context. We evaluate the performance both for generating the last image on the VIST dataset [34] and for generating multiple images auto-regressively on the Pororo [51] and Flintstones [28] datasets. Our model achieves SOTA performance even compared with previous specialist methods. Qualitative results can be found in Appendix.
4.3 Ablation Study
For ablation study, we use CLIP-ViT-L/14 [71] as the image encoder, OpenLLaMA-3B v2 [26] as the LLM and miniSD [^1] https://huggingface.co/justinpinkney/miniSD as the image decoder. The output image resolution is 256. The model is pre-trained for 10000 steps, using a mixture of LAION-COCO [78], LAION-2B [77], and MMC4 [111] datasets.
For simplicity, MMFS in this subsection is by default implemented in a single-image and single-scale setting, where each token only attends to the $16\times 16$ feature map of the nearest preceding image. We evaluate the zero-shot performance on three representative tasks and four datasets, i.e., image captioning on the COCO Karpathy test set, text-to-image generation on the COCO Karpathy test set, visual question answering on OKVQA [60], and TextVQA [81] validation set. The evaluation metrics are CIDEr, FID-5k, and top-1 accuracy, respectively. To ablate the cross-attention module in image generation, we also evaluate the fine-tuned results for segmentation-to-image translation task on ADE20K [108].